PS-SPS Users Meeting for Week 13 held on March 27th 2025

Coordinator for this week: Martin Jäkel

EuroLabs : please urgently submit a list of your publications to Tanya if you profited from EuroLabs funding.

News from the PS & SPS Physics Coordinator

M.R. Jäkel, P. Martinengo

Proton Run 2025

Newly vacant beam slots :

Please contact SPS.Physics.Coordinator@cern.ch if you want to apply for any of theses slots.

Reminder: Beam time exceeding the limits of 2 weeks PS beam time and 1 week SPS beam time per year need the approval of one of these CERN committees: SPSC, LHCC, DRDC, INTC, RB or IEFC. Consider joining a DRD collaboration, if you require more beam time.

Target intensities:
See below or via ASM (accessible from within the CERN network).

News from the Facilities Operations Meeting (FOM)

Upcoming SPS MDs next weeks :

https://be-dep-ea.web.cern.ch/content/md-planning-north-area
Dedicated SPS MDs - if not anounced differently - are taking place from 8:00 to 18:00


PS Machine Report (Y. Dutheil)

Safety (Marcus Jankowski)

n_ToF (Alice Manna and Paolo M Milazzo)

Three measurements are running:
• EAR1: Annular Silicon detector for (n, cp) reactions
• EAR2: Ionization chamber for fission studies (235U sample)
• NEAR: Study of radiation hardness for diamond detector and cycling proof of concept for activation measurements (see pictures).

EAR2: change of experimental set-up on 27.03: array of LaBr3 and an HP Germanium detectors for the study of the 87Sr(n, g) reaction

East Area Beam Status (Laurie Nevay)

On call phone number: 67500 (this week Johannes)

T09: Good operation week. The longer spill of 0.8 s with 15 GeV primary protons is being checked for the user. Should be ready to use towards the end of the week.
T10: Issue with beam stopper of T10, not letting the (hadron) beam be turned on for about 1 day. Some time was lost, some time was spent testing with muon beams. Final repair to be carried out during the next East Area beam stop, under discussion still.
T11: No user. Shutter closed until next operation.

East Area Users Tour de Table

T8 Main: IRRAD/CHARM (F. Ravotti, S. Fiore)

We restarted as soon as the East Area intervention on T10 was completed on Friday by performing a calibration of the IRRAD Beam Instrumentation with low intensity beams (E9-E10 pps). From Friday night we began the facilities commissioning with standard beam (8E11 pps) - this is mainly to calibrate the radiation field with various CHARM configurations. Access on Monday, Wednesday and Friday this week. In IRRAD we installed a first active experiment (SY-BI) on Wednesday, long-term material samples from EP-DT and TE-MSC will go in on Friday. First CHARM user teams will begin on Wednesday next week. We continue (announcing the daily plan using the operation e-group) until then.

T8 beam steering is in general very good (Thanks!). Please, watch out to certain spills (depending on CPS) that might be off on the X-plane of several mm.

Current list of registered IRRAD experiments so far (15 in total):

  1. 1x SY-BI (straws)
  2. 3x DRD3 and EP-RD (LGAD and high-level fluence detector samples)
  3. 1x LHCb SciFi upgrade
  4. 2x CMS-PPS (LGAD and cooling system samples)
  5. 1x EP-DT (polimeric materials)
  6. 2x ALICE (CMOS/LGAD and FOCAL)
  7. 1x EN-EL (GaN transistors for LED lights)
  8. 1x SY-STI and HSE-RP (development of lower momentum p+ beams)
  9. 3x TE-MSC (magnets isolation materials, RT, inert-gas and cryogenic)

For CHARM we have up to now 37 requests from users from ATS and EP

T9 Main: WCTE (M. Hartz)

T10 Main: BI STRAW (Robert Garbrecht Larsen)

We are continuing measuring on our straw detector at different energies and intensities. In general beam availability has been good.

T10 Incoming Main: MEDIPIX (Week 14) (Benedikt Bergmann)

North Area Beam Status (Laurie Nevay)

Incoming Main HIRADMAT (Week 14) (Please Put Your Name Here)

North Area Users Tour de Table

M2 Incoming Parasitic: AMBER (Week 14)


AoB


Minutes by the respective speakers, edited by E. B. Holzer, M. Jäkel, T. Shulha, and M. Schwinzerl